OpenStudy (anonymous):

Find the distance between (1, 2) and (4, 6). Round the answer to the nearest hundredth, if necessary

OpenStudy (anonymous):

Sqr rt of (delta x^2 + delta x^2)

OpenStudy (anonymous):

Sqr rt of [3^2 +4^2]=5

OpenStudy (anonymous):

To find the distance between two points on a graph, you should always create a right triangle out of it: (1,2) and (4,6) have a similar point (1,6). If you draw it out on a graph, it is easier to see because each point given follows either the x value or y value to get to this point. Now you can find the distance between the two given coordinates and the similar point: (1,6) is 4 units away from (1,2), so 4 is one leg of the right triangle. (1,6) is 3 units away from (4,6), so 3 is the other leg of the triangle. At this point, use the pythagorean theorem and plug in the two legs for A and B: a^2+b^2=c^2 3^2+4^2=c^2 9+16=c^2 25=c^2 Square root of 25 is 5, which is the distance between the two points. I hope you read over this and understand it instead of finding the answer among this work and putting it in.

OpenStudy (anonymous):

Thank you both !